Shift report 20000726 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Jul 26 2000 - 08:58:36 EDT

  • Next message: http account: "Shift report 20000726 08:00-16:00"

    ------------------------------------------------------------
    Date: Wed Jul 26 08:58:36 2000, Shift: 00:00-08:00
    Supervisor: PC, with: PS
    Log book#11 pages 81 - 92
    
    Shift Summary:
    ------------------------------------------------------------
    Beam at 70GeV from 1 am to 7 am. Run 1941. Looking back, we 
    should  probably have split it up in smaller bits. 
    They steered to us at ~01:40 and thereafter we had good beam
    for a while. At 3:30 they had steered to all experiments and
    beam was stable, but with few collisions.
    There was 1282 good collisisns out of 200K.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Jul 26 2000 - 09:01:55 EDT